OpenStudy (anonymous):

length of movie projected at 24 frames per second L=24f. solve for F?

OpenStudy (anonymous):

Length of movie projected at 24 frames per second: l = 24f solve for f - f=l/24 - 24xl=f because you have to get f by itself. - f = l/24

OpenStudy (anonymous):

give me a bage :)